OpenStudy (andijo76):

i am starting a new class and need help substitution method t=4-2s,t+2s=6

OpenStudy (the_fizicx99):

I don't understand your equation, I tried going by it 2 different ways and it didn't make sense..

OpenStudy (cloverracer):

retype your equations please. (:

OpenStudy (the_fizicx99):

I got -2 somehow. Could you specify your equation?

OpenStudy (agent0smith):

t=4-2s t+2s=6 ^ from the first equation, you know that t is (4-2s), so replace the t in t+2s=6 with (4-2s): (4-2s)+2s=6 now solve that for s

OpenStudy (andijo76):

what it said in my chapter is that for excersises 11 through 58 if a system has an infinate number of solutions use set builder notation to write the solution set if a system has no solution state this the substitution method solving using the substitution method i dont know if this helps

OpenStudy (the_fizicx99):

It doesn't have a solution, solve for s. and you'll see

OpenStudy (andijo76):

ok i will try lol (algebra is not my strong point )

OpenStudy (the_fizicx99):

Its parallel, and it doesn't have a solution therefore its false.

OpenStudy (andijo76):

sot that is the answer there is no solution that its false?

OpenStudy (the_fizicx99):

Yes, its false, unless its asking you for the answer then solve for s

OpenStudy (andijo76):

ok thankyou

OpenStudy (ranga):

Your instruction says: "if a system has no solution state this" This set of equations has no solutions. (After rearranging, t + 2s = 4, t + 2s = 6 are the two equations. This can only happen if 4 equals 6 because the left sides are the same. So this system has no solutions.)

OpenStudy (the_fizicx99):

@andijo76 Your welcome :)